A piece of alloy weights 0.86N in air and 0.73N When Immersed in water. Find volume and the density of the alloy


1
Expert's answer
2022-02-28T13:40:56-0500

The weight of alloy in water

"W=mg-g\rho V"

The volume of alloy

"V=\frac{mg-W}{g\rho}\\
=\frac{0.86-0.73}{9.8*1000}=1.33*10^{-5}\:\rm m^3"

The density of alloy

"\rho_a=\frac{m}{V}\\
=\frac{0.86}{9.8*1.33*10^{-5}}=6615\:\rm kg/m^3"